OpenStudy (anonymous):

factor 36x^2 + 24x -24

OpenStudy (mathmate):

36x^2 + 24x -24 =12(3x^2+2x-2)

OpenStudy (anonymous):

So then whats the factor of 3x^2 +2x -2?

OpenStudy (mathmate):

There are no rational factors.

OpenStudy (anonymous):

What are the irrational factors? If any

OpenStudy (mathmate):

\[ (x+(\sqrt(7)+1)/3)(3x-\sqrt(7)+1)\]

OpenStudy (paxpolaris):

If you are asked to factor 6 you would say 2*3, not sqrt(3)*sqrt(12) =12(3x^2+2x-2) I think this should be an acceptable final solution

OpenStudy (mathmate):

I think he understood that, but for curiosity, he wanted to know what the irrational factors are.